# A Survey on Federated Analytics: Taxonomy, Enabling Techniques, Applications and Open Issues

### v1

@@ II Overview of Federated Analytics
In this section, we provide an overview of the emerging FA. In Section II-A, we describe the general workflow of FA. In Section II-B, we propose a general mathematical problem formulation of FA. In Section II-C, we introduce the relationship between FA and related fields.
II-A General workflow
@@
FA aims at conducting data analytics tasks based on the federated data held by multiple clients with privacy preservation. The fundamental principle of FA to realize privacy preservation is that the raw data held by the clients are not transmitted and exposed. In practice, the prevention of raw data exposure is not sufficient to realize a formal privacy guarantee, and extra privatization mechanisms are usually applied in FA for the formal privacy guarantee. As is demonstrated in Fig. 2, the typical architecture of FA (server-client model) includes a server, which is hosting a data analytics tasks, and multiple clients, which would like to contribute their personal data to the data analytics task but require privacy preservation on their data. The right part of Fig. 2 demonstrates the steps of one-shot and iterative FA systems. In the FA workflow, the clients are in charge of a part of the data analytics task. They receive the computation model from the server, and perform local computation based on the received computation model and the client’s local data. The result of local computation, which is termed as “insight” in this survey, is uploaded to the server, where the insights are processed by privatization mechanisms to realize privacy preservation. The server aggregates the received insights with a dedicated aggregation algorithm. The aggregation output may be directly the results of the data analytics task when the FA algorithm is one-shot. When the algorithm is iterative, the aggregation result updates the computation model, and the data analytics results are derived after many rounds of model distribution, local computation, insight upload, and insight aggregation.
II-B Mathematical definition
@@
FA considers a system of $n$ data owners, or clients, denoted $DO$, and each client $DO_{i}$ holds local client data $x_{i}$. These clients would like to collaboratively accomplish a data analytics task, which can be reformulated as deriving the value of the task function
@@
Based on the aforementioned mathematical formulation, the research problem in designing the FA algorithm is derived: the researchers need to design insight derivation function $\mathcal{I}$ and aggregation function $\mathcal{A}$, so that the FA result $\mathcal{A}(\{\mathcal{I}(x_{i});i\in DO\})$ should be identical (or close to) the result of the task function $\mathcal{F}(x_{1},...,x_{n})$.
@@
II-C Comparison with related topics
TABLE I: Difference between FL and FA
@@ II-C Comparison with related topics
TABLE II: Difference between DDM and FA
DDM FA
@@
Although the term FA has been introduced in recent years, some core characteristics of FA, such as privacy protection, distributed computing, can already be found in existing research fields, such as FL, privacy-preserving data mining, and distributed data analytics. These topics naturally have some overlap with FA. We list these three related topics of FA, introduce the overlapping research, and discuss how FA extends their concepts.
II-C1 Relationship between FL and FA
@@
Both Fl and FA target at conducting intelligent data-oriented tasks without centralizing the raw data. Their major difference can be summarized in Table. II. First, FA and FL are differentiated for their goal tasks. In FL, the mechanism is designed to serve a predictive model training task. In most of the FL literature, the “model training” in FL exactly refers to deep supervised learning on a neural network, if no specific context is present. On the contrary, as its name refers, FA tackles the extensive field of descriptive data analysis tasks, ranging from the simplest task of calculating the average to complex tasks such as graph analytics and video analytics. Second, the task range of FA diversify the forms of local insights and central aggregation. In contrast to the model weights or gradient distilled from the local clients, the insights from FA could be much more diverse and task dependent, such as privatized hashing results [15], one-bit response to server query [16], or local clustering centroids [17]. The procedure of insight aggregation of FA is also diverse and task dependent compared to FL, such as tree aggregation [18], Bayesian-based distribution estimation [19], or secret sharing decryption [20].
II-C2 Relationship between FA and privacy-preserving data mining
@@
Privacy-preserving data mining is a comprehensive term including all data analytics schemes that take privacy concerns into consideration. Therefore, FA can be considered as a subfield of privacy-preserving data mining. As the need to perform data analytics with privacy preservation naturally exists, many solutions for privacy-preserving data analytics existed even earlier than the term federated computation (FL was first proposed in 2016 and FA was in 2020). These works (represented by RAPPOR [15]) do not use the term “federated”, but follow the similar idea and methodology as FA. Therefore, it is reasonable and scientific to re-classify these works as FA. On the other hand, not all privacy-preserving data mining research can be considered as FA. FA centers around distributed data setting and local privacy preservation. Heterogeneity and the resulting issues born at the client side, such as client dynamics, adversarial attacks, communication overhead, and incentive issues, are all new problems that are not considered in classical privacy-preserving data mining.
II-C3 Relationship between FA and distributed data mining
@@
Distributed data mining (DDM) refers to the procedure of conducting data analytics tasks while the data is distributed among multiple parties or machines. According to such a comprehensive definition, FA is a subfield of DDM because of its federated data setting. However, the FA studies differ from the general distributed data mining in that distributed data mining usually operates in trusted database scenarios with distributed machines supporting computing acceleration. In summary, in DDM, the computation nodes are naturally trusted, and the raw data transmission is allowed, which is not allowed in FA. Table II concludes the differences between FA and DDM.
@@
TABLE III: Examples of FA algorithms and systems following the taxonomy of this survey.
Reference Data analytics task Client scale Num. Iterations Coord. model Threat model
[21, 22, 23, 24] Mean computation $>$1,000 One-shot Server-client Model of LDP
[25, 26, 27, 28, 29] Mean computation $>$1,000 One-shot Server-client Model of cryptography
[30] Median/percentile computation 2 Iterative Decentralized Model of cryptography
[31] Median/percentile computation 200 One-shot Server-client Model of cryptography
[32] Median/percentile computation 15 Iterative Decentralized Hiding raw data
@@
[17] Clustering 200 One-shot Server-client Hiding raw data
[33, 34, 35, 36] Clustering 4$\sim$100 Iterative Server-client Hiding raw data
[15, 37, 38, 39, 40] Frequency oracle/Frequent pattern mining $>$1,000 One-shot Server-client Model of LDP
[41, 42, 16] Frequency oracle/Frequent pattern mining $>$1,000 Iterative Server-client Model of LDP
[18, 43] Frequency oracle/Frequent pattern mining $>$1,000 Iterative Server-client Model of CDP
[44] Frequency oracle/Frequent pattern mining $>$1,000 One-shot Server-client Models of cryptography & CDP
[45] Frequency oracle/Frequent pattern mining $>$1,000 Iterative Server-client Models of cryptography & CDP
[46, 47, 48] Query 2$\sim$3 One-shot Decentralized Model of cryptography
[49] Query 2 One-shot Decentralized Models of cryptography & CDP
@@
[50, 51, 52] Query 10$\sim$16 One-shot Server-client Model of cryptography
@@
[53, 54, 55] Query $>$1000 Iterative Server-client Model of cryptography & CDP
[56] Matrix factorization 610 One-shot Server-client Model of cryptography
[57] Matrix factorization 40 Iterative Server-client Model of cryptography
[58] Matrix factorization 2 One-shot Decentralized Model of cryptography
@@
[59, 60] Multi-armed bandit 10 Iterative Server-client Hiding raw data
[61, 62, 63] Multi-armed bandit 10 $\sim$100 Iterative Decentralized Model of CDP
III Taxonomy of Federated Analytics
@@
Given the diversity of the existing FA algorithms and systems, it becomes beneficial to provide methods to classify FA problems and solutions. In this section, we provide the taxonomy of FA. It examines the design and application of FA from five important dimensions, including data analytics tasks, scale of clients, number of iterations, coordination models, and threat models.
III-A Data analytics tasks
@@
As the application of FA covers massive and diverse data analytics needs, classifying FA algorithms and systems based on the types of data analytics tasks they solve becomes a natural and useful taxonomy of FA. The host data analytics task heavily influences the algorithmic design, characteristics, and challenges of the corresponding FA algorithms and systems, and the algorithmic design of FA is expected to be quite different when handling different types of data analytics tasks. One of its reasons is that different data analytics tasks extract and analyze different features form the raw data, which then determine the structure of required insight, the insight generation approach, and insight aggregation approach. Some researcher would like to break the limit of data analytics task-specific FA design, and propose some unified FA framework. However, proposing such an FA framework that can handle all data analytics tasks is still far from reality at this moment. Therefore, the host data analytics task is currently an important component of any FA taxonomy. In Section XII-C, we investigate the probability of a unified FA framework as a future research direction.
In this survey, we follow the taxonomy of data analytics tasks to introduce the existing FA research. We discuss five classes of data analytics applications, in Sections VI-X respectively, and introduce the existing FA algorithms and systems in the context of the host data analytics tasks.
III-B Scale of clients
FA algorithms and systems have diverse settings regarding the scale of participating clients, from fewer than 10 to more than a million. The difference has a large relationship with the privacy mechanisms they use as well as their application scenarios. For mobile applications, an FA scheme is usually expected to maximize its capacity to support as many participating clients as possible. For industrial scenarios with only a few data silos, an FA scheme is also expected to work, usually with higher requirements on analytics performance.
@@ III-B Scale of clients
Usually, DP-based mechanisms are better at supporting massive participating clients, because they only require the clients to individually perturb their client uploads in parallel, and the server aggregation function usually has linear complexity [21, 22, 15, 37, 16]. In addition, when the privacy model of CDP is considered, increasing the number of participating clients can increase privacy preservation by better hiding the information of individual clients [18, 43].
@@
On the other hand, FA schemes utilizing computation-intensive cryptography tools, though can provide higher analytics utility, are usually hindered by the algorithm, computation, or communication limits. Consequently, they can only be applied to a smaller scale of clients. For example, the garbled circuit [64], a classical and famous cryptography tool employed in some FA studies [46, 50], only supports computation between two parties, and modifications of it usually introduces significant computation and communication overhead to support several more participating clients. To overcome the limit in supported client scale when utilizing cryptography tools, many research efforts are proposed to increase the supported participating clients in cryptography-based FA systems. An example direction of these research is to replace the computation-intensive cryptography tools by lightweight ones, like simple additive masking [25, 26, 27, 28].
III-C Number of iterations
FA algorithms can be classified into one-shot (one-round) non-iterative schemes and iterative schemes. The one-shot schemes are completed after conducting model distribution, local computation, and insight aggregation once. In contrast, the iterative schemes repeat the phases, constantly deriving new knowledge and optimizing the computation model, until the data analytics task is accomplished. Compared to the one-shot algorithms, iterative algorithms usually require more complex design, because iterative algorithms requires extra considerations about how to maintain the “state” of the data analytics procedure, and how the local computation procedure of the clients can be guided by the global knowledge derived in previous rounds. Roughly speaking, one-shot schemes are usually applicable in simple FA tasks and incur less communication cost, like federated mean computation [21, 22], while iterative schemes are usually required for complex FA tasks with explicit computation models, like federated multi-armed bandits [59, 60, 61, 62, 63].
III-D Coordination models
The vanilla setting of FA considers the server-client architecture, where all the clients only communicate with the centralized server. The server-client architecture becomes the mainstream form of FA because the clients can easily register themselves in the system, the FA system can easily scale up, and the communication scheme can be easily determined. However, as some shortcomings of the server-client architecture are observed by the researchers, alternative architectures, such as server-client architecture with client-client communication [25, 26], and even a fully decentralized architecture without any centralized server [61, 62, 63], are proposed. These alternatives have potential advantages such as preventing one-point failure, better privacy preservation, better communication efficiency, and adaptive organization.
III-E Threat models
While FA algorithms and systems apply different privatization techniques to guarantee privacy preservation (Section V-A provides a demonstration), different threat models are considered that correspond to these privatization techniques. For example, in these FA systems utilizing cryptography tools like homomorphic encryption or multi-party computation, the threat model can be assumed that privacy is preserved when the server cannot know the decrypted upload from the clients; in these LDP-based FA schemes, the threat model prevents the adversary from inferring the client raw data by analyzing the client upload; in these CDP-based FA schemes, the threat model prevents the adversary from inferring the client raw data by analyzing the final analytics results. There does not exist an inclusion relationship between the threat model of cryptography tools and the threat model of DP, and one cannot claim that one is strictly stronger than the other. In addition, FA threat models can also be classified regarding the trust in the server or aggregator. Few works assume that there exists a trusted server or aggregator; the majority of FA works assume that the server is honest-but-curious (also known as semi-honest); the rest of FA research investigates privacy reservation over stronger adversarial models.

### v3

@@ III Overview of Federated Analytics
In this section, we provide an overview of the emerging FA. We focus on many important concepts of FA. In Section III-A, we describe the general workflow of FA. In Section III-B, we propose a general mathematical problem formulation of FA. In Section III-C, we introduce the relationship between FA and related fields. In Section III-D, we present taxonomies of FA in five dimensions that help us properly position the diverse FA algorithms and systems.
III-A General workflow
@@
FA aims at conducting data analytics tasks based on the federated data held by multiple clients with privacy preservation. The fundamental principle of FA to realize privacy preservation is that the raw data held by the clients are not transmitted and exposed. In practice, the prevention of raw data exposure is not sufficient to realize a formal privacy guarantee, and extra privatization mechanisms are usually applied in FA for the formal privacy guarantee. Since FA and FL are two variations of the federated computation paradigm, they share identical abstractive architecture and workflow, which is defined in the federated computation paradigm. As is demonstrated in Fig. 2111The right plot of the figure only applies for iterative FA schemes according to our taxonomy of iteration pattern (Section III-D)., the typical architecture of FA (server-client model) includes a server, which is hosting data analytics tasks, and multiple clients, that would like to contribute their data to the data analytics task but requires privacy preservation on them.
The right part of Fig. 2 demonstrates the steps of one-shot and iterative FA systems. In the FA workflow, the clients are in charge of a part of the data analytics task. They receive the computation model from the server and perform local computation based on the received computation model and the client’s local data. The result of local computation, which is termed “insight” in this survey, is uploaded to the server, where the insights are processed by privatization mechanisms to realize privacy preservation. The server aggregates the received insights with a dedicated aggregation algorithm. The aggregation output may be directly the results of the data analytics task when the FA algorithm is one-shot. When the algorithm is iterative, the aggregation result updates the computation model, and the data analytics results are derived after many rounds of model distribution, local computation, insight upload, and insight aggregation.
As mentioned above, the same architecture and workflow of FA is also adopted by many classical FL systems. However, it worth to be emphasizing that FA and FL differ a lot when the abstractive architecture and workflow are implemented into specific algorithms and programs. FL and FA perform different computations on the client side, deriving different structures of insights, and are eventually aggregated by the server in different ways. The difference in downstream tasks also affects the abstractive workflow of FA and FL. For example, one-shot schemes (e.g., those with only iterations) are much more common in FA than FL, because they are suitable for deriving simple statistics from client data. More discussions and comparisons are present in the following Section III-C.
III-B Mathematical definition
@@
FA considers a system of $n$ data owners, or clients denoted $DO$, and each client $DO_{i}$ holds local client data $x_{i}$. These clients would like to collaboratively accomplish a data analytics task, which can be reformulated as deriving the task function
@@
Based on the aforementioned mathematical formulation, the research problem in designing the FA algorithm is derived: the researchers need to design insight derivation function $\mathcal{I}$ and aggregation function $\mathcal{A}$ so that the FA result $\mathcal{A}(\{\mathcal{I}(x_{i});i\in DO\})$ should be identical (or close to) the result of the task function $\mathcal{F}(x_{1},...,x_{n})$.
@@
III-C Comparison with related topics
TABLE II: Difference between FL and FA
@@ III-C Comparison with related topics
TABLE III: Difference between non-FA parts of DDM and FA
non-FA DDM FA
@@
Although the term FA has been introduced in recent years, some core characteristics of FA, such as privacy protection, and distributed computing, can already be found in existing research fields, such as FL, privacy-preserving data mining, and distributed data analytics. These topics naturally have certain overlap with FA. We list these three related topics of FA, introduce the overlapping research, and discuss how FA extends its concepts. We also provide illustration of their relationships in Fig. 3, where FA lies in the overlap region of distributed data mining, privacy-preserving data mining, and federated computation, while FL, which focus on neural network training rather than data mining, is another variation of the federated computation paradigm.
Figure 3: Relationship of FA and related topics.
III-C1 Relationship between FL and FA
@@
Both FL and FA target at conducting intelligent data-oriented tasks without centralizing the raw data. Their major difference can be summarized in Table III. First, FA and FL are differentiated for their goal tasks. In FL, the mechanism is designed to serve a predictive model training task. In most of the FL literature, the “model training” in FL exactly refers to deep supervised learning tasks, if no specific context is present. These FL algorithms utilize gradient descent (typically on a neural network). On the contrary, as its name refers, FA tackles the extensive field of descriptive data analysis tasks, ranging from the simplest task of calculating the average to complex tasks such as graph analytics and video analytics [23, 24]. Second, the task range of FA diversifies the forms of local insights and central aggregation. In contrast to the model weights or gradient distilled from the local clients, the insights from FA could be much more diverse and task-dependent, such as privatized hashing results [25], one-bit response to server query [26], or local clustering centroids [27]. The procedure of insight aggregation of FA is also diverse and task-dependent compared to FL, such as tree aggregation [28], Bayesian-based distribution estimation [29], or secret sharing decryption [30]. To be specific, in this survey, if a federated computation scheme is based on an explicit parameterized computation model, and computes by gradient-based methods on the parameterized model, such a scheme will be considered as FL. A federated computation scheme is considered FA rather than FL when it conducts traditional data analytics task or performs computation by a non-gradient descent method.
III-C2 Relationship between FA and privacy-preserving data mining
@@
Privacy-preserving data mining is a comprehensive term including all data analytics schemes that take privacy concerns into consideration. Therefore, FA can be considered as a subfield of privacy-preserving data mining. As the need to perform data analytics with privacy preservation naturally exists, many solutions for privacy-preserving data analytics existed even earlier than the term federated computation (FL was first proposed in 2016 and FA was in 2020). These works (represented by RAPPOR [25]) do not use the term “federated”, but follow a similar idea and methodology as FA. Therefore, it is reasonable and scientific to re-classify these works as FA. On the other hand, not all privacy-preserving data mining research can be considered as FA. FA centers around distributed data setting and local privacy preservation. Heterogeneity and the resulting issues born on the client side, such as client dynamics, adversarial attacks, communication overhead, and incentive issues, are all new problems that are not considered in classical privacy-preserving data mining.
III-C3 Relationship between FA and distributed data mining
@@
Distributed data mining (DDM) refers to the procedure of conducting data analytics tasks while the data is distributed among multiple parties or machines. According to such a comprehensive definition, FA is a subfield of DDM because of its federated data setting. However, the FA studies differ from general distributed data mining in that distributed data mining usually operates in trusted database scenarios with distributed machines supporting computing acceleration. In summary, in conventional DDM, the computation nodes are naturally trusted, and the raw data transmission is allowed, which is not allowed in FA. Table III concludes the differences between FA and DDM.
Figure 4: An illustration of our taxonomy.
@@
TABLE IV: Examples of FA algorithms and systems following the taxonomy of this survey. The ubiquitous privatization methodology of hiding raw data is omitted in the rows if any other methodology is present.
Reference Data analytics task Data owner type Iter. pattern Coord. model Privatization methodology
[31, 32, 33, 34] Statistical metrics cross-device One-shot Server-client Local differential privacy
@@
[24] Statistical metrics cross-device One-shot Server-client Hiding raw data
@@
[35, 36, 37] Statistical metrics cross-device One-shot Server-client Cryptography
[38, 39] Statistical metrics cross-device One-shot Hybrid Cryptography
[40] Statistical metrics cross-silo Iterative Decentralized Cryptography
[41] Statistical metrics cross-device One-shot Server-client Cryptography
[42, 43, 44] Statistical metrics cross-device Iterative Server-client Central differential privacy
[45, 46, 47, 48, 49, 50] Statistical metrics cross-device Iterative Server-client Local differential privacy
[51] Statistical metrics cross-device Iterative Decentralized Cryptography & Local differential privacy
@@
[52, 53] Statistical metrics cross-device Iterative Server-client Cryptography
[30] Statistical metrics cross-device Iterative Decentralized Cryptography & $k$-anonymity
[54] Statistical metrics cross-device Iterative Server-client Cryptography
[55] Statistical metrics cross-device One-shot Decentralized Central differential privacy & $k$-anonymity
[56] Statistical metrics cross-silo Iterative Decentralized Hiding raw data
[27] Statistical metrics cross-device One-shot Server-client Hiding raw data
@@
[57] Statistical metrics cross-device Iterative Server-client Hiding raw data
[58, 59, 60] Statistical metrics cross-silo Iterative Server-client Hiding raw data
[25, 61, 62, 63, 64, 65] Frequency-based applications cross-device One-shot Server-client Local differential privacy
[66, 26, 67, 68, 69] Frequency-based applications cross-device Iterative Server-client Local differential privacy
[28, 70] Frequency-based applications cross-device Iterative Server-client Central differential privacy
[71] Frequency-based applications cross-device One-shot Server-client Cryptography & Central differential privacy
[72, 17] Frequency-based applications cross-device Iterative Hybrid Distributed differential privacy
[73] Frequency-based applications cross-device Iterative Server-client Cryptography & Local differential privacy
[74, 75, 76] Database operations cross-silo One-shot Decentralized Cryptography
[77] Database operations cross-silo One-shot Decentralized Cryptography & Central differential privacy
[78, 79, 80] Database operations cross-silo One-shot Server-client Cryptography
[81, 82, 83] Database operations cross-device Iterative Server-client Cryptography & Central differential privacy
[15, 84, 85, 86, 87] Assisting federated learning cross-device Iterative Server-client Hiding raw data
[88] Assisting federated learning cross-device One-shot Server-client Hiding raw data
[89, 90, 91, 92, 14] Assisting federated learning cross-device Iterative Server-client Hiding raw data
[93, 94] Wireless network application cross-silo Iterative Server-client Hiding raw data
[95, 96, 97, 98, 99] Wireless network application cross-device Iterative Server-client Hiding raw data
III-D Taxonomy of federated analytics
@@
Given the diversity of the existing FA algorithms and systems, it becomes beneficial to provide methods to classify FA problems and solutions. In this section, we provide the taxonomy of FA. It examines the design and application of FA from five important dimensions, including data analytics tasks, scale of clients, number of iterations, coordination models, and threat models. Fig. 4 illustrates the taxonomy.
III-D1 Data analytics task
As the application of FA covers massive and diverse data analytics needs, classifying FA algorithms and systems based on the types of data analytics tasks they solve becomes a natural and useful taxonomy of FA. In this survey, we categorize the corresponding data analytics tasks of the FA works we surveyed into five categories: statistical metrics, frequency-related applications, database operations, assisting FL applications, and wireless network applications. Following the taxonomy, we respectively introduce FA works focusing on each data analytics task in Sections VI-X. In some sections, a more detailed classification of the data analytics tasks is present.
@@ III-D1 Data analytics task
The host data analytics task heavily influences the algorithmic design, characteristics, and challenges of the corresponding FA algorithms and systems, and the algorithmic design of FA is expected to be quite different when handling different types of data analytics tasks. One of its reasons is that different data analytics tasks extract and analyze different features from the raw data, which then determine the structure of required insight, the insight generation approach, and the insight aggregation approach. Some researchers would like to break the limit of data analytics task-specific FA design, and propose some unified FA framework [26]. However, proposing such an FA framework that can handle all data analytics tasks is still far from reality at this moment. Therefore, the host data analytics task is currently an important component of any FA taxonomy. In Section XI-B, we investigate the probability of a unified FA framework as a future research direction.
III-D2 Data owner type
FA algorithms and systems can be categorized into cross-device FA and cross-silo FA based on the essence of their participating clients (data owners). Cross-device FA usually involves massive (millions of) clients with little data and limited computation/communication capacity, while cross-silo FA usually involves a small number of clients, with more data and stronger resources, which have clear objectives for participating in the federation. Cross-device FA and cross-silo FA have significant differences in many aspects, including the utilized privacy mechanisms and application scenarios. For example, cross-device FA is usually deployed in mobile applications with massive edge devices. Therefore, the cross-device FA mechanisms are required to easily scale up, and introduce limited workload to the clients. On the other hand, for cross-silo FA in industrial scenarios (e.g., finance, biology), there might be only a few data silos, and each of them possesses a large number of high-quality data. In these cases, the participants are likely to derive highly correctness analytics results, introducing little error compared to the centralized algorithms. At least the performance of FA should be higher than centralized analytics on its local database, i.e., the benefit of collaboration should at least cover the degrading in privacy preservation.
@@ III-D2 Data owner type
Usually, DP-based mechanisms are better at supporting cross-device FA, because they only require the clients to individually perturb their client uploads in parallel, and the server aggregation function usually has linear complexity [31, 32, 25, 61]. In addition, when the privacy model of central differential privacy is considered, increasing the number of participating clients can increase privacy preservation by better hiding the information of individual clients [28, 70].
@@
On the other hand, FA schemes utilizing computation-intensive cryptography tools, though can provide higher analytics utility, are usually hindered by the algorithm, computation, or communication limits. Consequently, they are usually applied to cross-silo FA. For example, the garbled circuit [100], a classical and famous cryptography tool employed in some FA studies [74, 78], only supports computation between two parties, and modifications of it usually introduces significant computation and communication overhead to support several more participating clients. To overcome the limit in supported client scale when utilizing cryptography tools, many research efforts are proposed to increase the supported participating clients in cryptography-based FA systems. An example direction of this research is to replace the computation-intensive cryptography tools with lightweight ones, like simple additive masking [38, 39, 35, 36].
III-D3 Iteration pattern
FA algorithm involves steps of computation model distribution, local computation, insight upload, and insight aggregation, where the aforementioned steps compose one FA iteration. FA algorithms can then be classified into one-shot FA and iterative FA, based on their patterns in conducting iterations.
In one-shot algorithms, the analytics results are derived and the algorithm terminates after one iteration. Since the system only processes one round, there is no memorized “state” on the server side, and the computation model of the FA algorithm is fixed. Therefore, the phase of computation model distribution can even conducted along with the deployment of the FA program. Roughly speaking, one-shot schemes are usually applicable in simple FA tasks and incur less communication cost, like federated mean computation [31, 32].
On the contrary, other FA algorithms involve multiple iterations until completion. In each iteration, a part of the analytics tasks is conducted, which updates the “state” on the server side. Based on the results in previous rounds, the server then provides information to participating clients to guide their local computation. There exist two fundamental ideas (ways) to design an iterative FA algorithm. Firstly, each iteration of FA aims to conduct different parts of the analytics task, where the computation of the later part usually depends on the results of the former parts. For example, many FA solutions for frequent pattern mining [28, 70, 67] let different FA iterations reveal frequent sequential patterns with different fixed lengths, where the computation on longer length depends on the shorter ones. In the second way, the computation model is formulated as the optimization of a parameterized model. In each round, the uploaded insight helps the optimization of the model (e.g., the insight is essentially the gradients on the model, being consistent with the methodology of FL). Unlike the former category, FA algorithms in the second category can also choose to terminate when the model converges instead of terminating after a predefined fixed number of iterations. Examples of the second category include clustering solutions [59] and assisting FL solutions [92, 87].
III-D4 Coordination model
In our previous introduction, we define the entities of FA as one server and multiple clients, where the server hosts the data analytics tasks by distributing the computation model, receiving insights, and aggregating insights, and the clients contribute their local data by conducting local computation and uploading insights. Such systems formulate the vanilla server-client architecture of FA, where all the clients only communicate with the centralized server. The server-client architecture has become the mainstream form of FA for multiple reasons. First, in the typical server-client architecture, the clients are unaware of other clients. Therefore, a new client can easily register itself in the FA system simply via communicating with the server. Secondly, an FA system under the server-client architecture can easily scale up, because the computation and communication complexity of the client scheme are usually constant, under such architecture. Thirdly, the complexity of the corresponding algorithm design is simplified, since the communication pattern has been defined by the architecture, and the privacy protection model is relatively simple.
However, as some shortcomings of the server-client architecture are identified by the researchers, optimized alternative architectures, are proposed. We classify them as hybrid architecture, where a server-client system is considered, but client-client communication is introduced in addition to the server-client communication. In [38, 39], the authors propose some FA solutions under hybrid architecture. By leveraging pairwise communication, the clients can enforce encryption on their uploads at a lower cost, and therefore collaborate to enhance privacy preservation against the potentially malicious server.
In [40, 51, 56], fully decentralized architectures without any centralized server are introduced. Fully decentralized FA algorithms are usually more complex in design but have the potential to obtain advantages including prevention of one-point failure, better privacy preservation, better communication efficiency, and adaptive organization.
III-D5 Privatization methodology
The tenet of FA, as well as more broadly federated computation, is to protect the data privacy of edge data while leveraging these data in data-intensive applications. Therefore, each FA work has to follow one or multiple methodologies in their algorithmic design. As a result, classifying the FA works based on the approaches to enforce privacy preservation becomes an appropriate and beneficial taxonomy in the field of FA. FA works following different privatization methodologies scatter different threat models, and techniques focuses. They can be applied to support diverse use requirements. This taxonomy is highly associated with various privatization techniques leveraged by FA, where Section V-A provides a thorough demonstration of these techniques. In this survey, we summarize six different privatization methodologies: hiding raw data, local differential privacy, central differential privacy, distributed differential privacy, $k$-anonymity, and cryptography.
The methodology of hiding raw data is the fundamental idea of federated computation, and it is applied in all FA works to provide privacy preservation to some extent. Under the privatization methodology of hiding raw data, the clients will extract insights from the raw data and upload the insights to the server for further data analytics. It prevents the extreme privacy risk in conventional data analytics where the server gathers all raw data from the clients. Since the extracted insight is an indirect transformation of the raw data where the reverse computation is usually not available, and the extracted insight usually possesses a much smaller dimension size than the original data, the server or other attackers cannot easily infer the raw data from the insights; The data privacy of the raw data is thus protected. However, there remain privacy risks as the attacker can still obtain valuable and sensitive information if the insights are naïvely extracted. Therefore, many FA solutions apply additional privatization methodologies using modern privacy preservation technologies (Section V-A) to further improve its privacy preservation, which we will discuss next.
DP-based privacy preservation methodologies, including local differential privacy (LDP), central differential privacy (CDP), and distributed differential privacy (DDP), are the most widely used additional privatization methodologies in the FA literature, since they provide rigid theoretical guarantees from the statistical perspective. Namely, the attacker cannot infer the existence of any raw data with sufficient confidence from the DP-possessed insights. LDP is the most popular one among DP methodologies because it assumes a local model that fits the settings of the majority of FA scenarios. In the methodology of LDP, the clients individually possess their local insights by adding perturbations to obtain sufficient privacy preservation. CDP considers protecting data from each client in the aggregated results. CDP requires a much smaller magnitude of perturbations, which results in higher data analytics quality but assumes the existence of a trusted aggregator who gathers raw data from the clients and is therefore not desirable in most FA scenarios. DDP is a novel idea of privatization methodology that combines the advantages of CDP and LDP by providing local privacy preservation with CDP-level perturbation magnitude, owing to its ingenious involvement of cryptography tools. DDP is gaining increasing research interest from FA researchers.
$k$-anonymity is another privatization methodology from the statistical perspective. It protects the privacy of client data (or its transformed form) by removing all the records with less than $k$ records. Therefore, a record can be seen by the server only when it has at least $k$ clients holding that record.
Cryptography is the last privatization methodology we focus on in this survey. Utilizing various tools from the field of cryptography, it provides privacy preservation from a different perspective from statistics. It focuses on encrypting the whole upload from all clients so that an adversary can only learn the aggregated results, but cannot learn any upload from individual clients. As the statistics perspective and cryptography perspective consider the elimination of different privacy risks, neither of them can thoroughly outperform the other. Some recent FA research even utilizes multiple privatization methodologies to enhance privacy preservation from multiple perspectives. The merge of cryptography and CDP also yields the emerging privatization methodology of DDP, which combines advantages and enhanced privacy preservation.